[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Mon Apr 17, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Mon Apr 17 18:47:01 UTC 2017


TAFB Shift Log for Mon Apr 17 2017 14:46:55 GMT-0400 (EDT)

Shift: Day Shift

Pacific Forecaster: S Stripling

Atlantic Forecaster: A Levine

Surface Analyst: M Formosa

Technical Issues Summary:

GFE/NCP issues: Mello did a reconfig on LX2 and LX3 and are now configured  
as it was before upgrade 2 months ago. This took 2 months to take care of.  
Thanks Chris!

N-AWIPS issues: MRF had issues with Cumulus today as it locked up for over  
20 minutes and had to do a hard reboot. SS has some minor problems with  
Grunt, stalling for 3-5 minutes 3 different times today, but it returned to  
normal ops.

Other issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170417/9d862522/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list